40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2664     Entry time: Tue Feb 23 13:43:57 2021
Author: aaron 
Type: DailyProgress 
Category: PSOMA 
Subject: condensed fiber layout 

enter Tue Feb 23 13:43:48 2021

  • photographed the AI chassis innards. Where is the mini hdmi cable for the camera?
    • When I turned on the AI chassis, the Sorensen hit a current limit on + supply. I didn't notice this at first, so haven't checked it out (could be the limit is set too low, as AI board was fine on the benchtop). For now, do not turn on the AI chassis.
    • Without the AI chassis, the current draw is 2.2 A on positive supply, 1.2 A on negative. I'll try to identify where the extra 1 A is going, and see if I have a photo from earlier this year. I did find this elog from Johannes that shows an extra 1A on positive supply. Some of the cryo cavs equipment is still on, so I am not overly concerned.
  • redid the fiber components (attachment). Didn't touch any free space mirrors, but rotated the launches slightly (axisymmetric about z, no pit/yaw/z changes) when unscrewing fiber.

Next planning to add some foam around the fibers and compare beat spectrum before/after.

exit Tue Feb 23 22:16:33 2021

Attachment 1: 2AAF19D0-3F40-4146-9975-8EF4545C91A7.jpeg  3.455 MB  Uploaded Thu Feb 25 15:03:27 2021  | Hide | Hide all
2AAF19D0-3F40-4146-9975-8EF4545C91A7.jpeg
ELOG V3.1.3-